### v1

# Homological Filling and Minimal Varifolds in Four-Dimensional Einstein Manifolds

Abstract.
We study the smallest area $A_{\min}(M,g)$ of a $2$-dimensional stationary integral varifold in a closed Einstein $4$-manifold $(M^{4},g)$ with
$\operatorname{Ric}_{g}=\lambda g,\quad|\lambda|\leq 3,\quad\operatorname{Vol}(M,g)\geq v>0,\quad\operatorname{diam}(M,g)\leq D,\quad H_{1}(M;\mathbb{Z})=0.$
Building on the previous work [23] on homological filling functions, we show that for every $(M^{4},g)$ in this Einstein class there is an upper bound
$A_{\min}(M,g)\leq F_{\operatorname{Ein}}(v,D),$
where $F_{\operatorname{Ein}}$ depends only on $(v,D)$ and on quantitative Sobolev and $\varepsilon$–regularity constants for Einstein metrics.
1. Introduction
Let $(M^{4},g)$ be a closed Riemannian manifold and denote by $A_{\min}(M,g)$ the area of a smallest $2$-dimensional stationary integral varifold in $(M,g)$. In general, the existence of such a minimal object follows from the Almgren–Pitts min–max theory [20], and in low codimension and low dimension one can often upgrade the varifold to a smooth embedded minimal hypersurface using the regularity theory of Schoen–Simon–Yau [22, 21].
In this article we are interested in bounding $A_{\min}(M,g)$ in terms of global geometric data in the special case when $(M,g)$ is an Einstein $4$-manifold subject to natural non-collapsing and diameter bounds. Throughout we assume that
(1) $\operatorname{Ric}_{g}=\lambda g,\quad|\lambda|\leq 3,\quad\operatorname{Vol}(M,g)\geq v>0,\quad\operatorname{diam}(M,g)\leq D,\quad H_{1}(M;\mathbb{Z})=0.$
We write $\mathcal{E}_{1}(4,v,D)$ for the class of all closed Einstein $4$-manifolds satisfying (1).
This paper is a sequel to our previous work [23], where we considered the larger class with $|{\rm Ric}_{g}|\leq 3,\ \operatorname{Vol}(M,g)\geq v>0,\ \operatorname{diam}(M,g)\leq D,\ H_{1}(M;\mathbb{Z})=0,$ and proved a linear upper bound for the first homological filling function $\operatorname{FH}_{1}$ of $(M,g)$. More precisely, for every $(M^{4},g)\in\mathcal{M}_{1}(4,v,D)$ we showed that there exist functions $f_{1}(v,D),f_{2}(v,D)>0$ such that
$\operatorname{FH}_{1}(l)\ \leq\ f_{1}(v,D)\,l+f_{2}(v,D)\qquad\text{for all }l\geq 0.$
Combined with an effective version of the Almgren–Pitts min–max theory due to Nabutovsky–Rotman [19], this yields
$A_{\min}(M,g)\ \leq\ F(v,D)$
for some function $F$ depending only on $v$ and $D$. In that setting, however, the dependence of $f_{1},f_{2}$ and $F$ on the geometric parameters was not made explicit, since the constants coming from the Cheeger–Naber bubble–tree decomposition [8] were only known to exist abstractly.
In the present work we revisit this question in the more rigid Einstein setting. For $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ the curvature solves an elliptic system, and global Sobolev inequalities can be used, following Anderson [2, 3], to derive quantitative $L^{2}$–curvature bounds and $\varepsilon$–regularity with constants depending only on a Sobolev constant. This analytic input, combined with the Einstein specialization of the Cheeger–Naber bubble–tree decomposition and an improved combinatorial filling estimate, allows us to obtain an explicit upper bound for $A_{\min}(M,g)$ in terms of $(v,D)$ and the Sobolev constants.
Our first main result is the following quantitative bound for minimal $2$-varifolds in Einstein $4$-manifolds.
Theorem 1.1.
For every $v,D>0$ there exists a constant $F_{\operatorname{Ein}}(v,D)>0$ with the following property. If $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$, then the area $A_{\min}(M,g)$ of a smallest $2$-dimensional stationary integral varifold in $(M,g)$ satisfies
$A_{\min}(M,g)\ \leq\ F_{\operatorname{Ein}}(v,D).$
Moreover $F_{\operatorname{Ein}}(v,D)$ can be written explicitly in terms of the Sobolev constant $C_{S}(v,D)$, the global $L^{2}$–curvature bound, and the quantitative Einstein $\varepsilon$–regularity constants.
As in [23], the proof of Theorem 1.1 proceeds via a homological filling inequality expressed in terms of the first homological filling function $\operatorname{FH}_{1}$ of $(M,g)$.
Theorem 1.2.
For every $v,D>0$ there exist functions
$f_{1}^{\operatorname{Ein}}(v,D),f_{2}^{\operatorname{Ein}}(v,D)>0,$
with the following property. If $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ and $C\in\mathcal{Z}_{1}(M;\mathbb{Z})$ is a singular Lipschitz $1$-cycle, then $C$ bounds a singular $2$-chain $E\in\mathcal{C}_{2}(M;\mathbb{Z})$ such that
$\operatorname{mass}_{2}(E)\ \leq\ f_{1}^{\operatorname{Ein}}(v,D)\,\operatorname{mass}_{1}(C)+f_{2}^{\operatorname{Ein}}(v,D).$
Equivalently, the first homological filling function of $(M,g)$ satisfies
$\operatorname{FH}_{1}(l)\ \leq\ f_{1}^{\operatorname{Ein}}(v,D)\,l+f_{2}^{\operatorname{Ein}}(v,D)\qquad\text{for all }l\geq 0.$
Theorem 1.1 is then an immediate consequence of Theorem 1.2 and the min–max theorem of Nabutovsky–Rotman [19], which relates $A_{\min}(M,g)$ to $\operatorname{FH}_{1}$. In the Einstein setting we can track all constants arising from the Sobolev inequality, the $L^{2}$–curvature bound and the bubble–tree decomposition, thus making $F_{\operatorname{Ein}}(v,D)$ completely explicit once the analytic data are fixed.
2. Sobolev inequality and Einstein $\varepsilon$-regularity
Throughout this section we assume $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$.
2.1. Global Sobolev inequality
We use the $L^{2}$-Sobolev inequality on $(M,g)$, which is equivalent to a uniform isoperimetric inequality under our assumptions.
Theorem 2.1 (Sobolev inequality).
There exists a constant $C_{S}(v,D)>0$ depending only on $v,D$ such that for every $u\in C^{\infty}(M)$ we have
(2) $\Bigl(\int_{M}|u|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{M}|\nabla u|^{2}+\frac{1}{D^{2}}\int_{M}u^{2}\right)^{1/2}.$
Proof.
Since $|\operatorname{Ric}_{g}|\leq 3$, we have $\operatorname{Ric}_{g}\geq-3g$. By Bishop–Gromov volume comparison, there exists a dimensional constant $c_{0}>0$ such that
$\operatorname{Vol}B_{1}(x)\ \geq\ c_{0}\,v\qquad\text{for all }x\in M.$
In particular $M$ is non-collapsed on the unit scale and has a uniform lower bound on the isoperimetric constant: there exists $I(v,D)>0$ such that for all smooth domains $\Omega\subset M$ with volume at most $\frac{1}{2}\operatorname{Vol}(M)$ one has
$\operatorname{Vol}(\Omega)^{3/4}\ \leq\ I(v,D)\,\mathcal{H}^{3}(\partial\Omega),$
where $\mathcal{H}^{3}$ denotes the $3$-dimensional Hausdorff measure induced by $g$.
By a standard argument (see e.g. Hebey, Nonlinear analysis on manifolds), such inequality is equivalent to an $L^{2}$-Sobolev inequality: there exists $C_{S}(v,D)$ such that
(3) $\Bigl(\int_{M}|u|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{M}|\nabla u|^{2}\right)^{1/2}$
for every $u\in C^{\infty}(M)$ with zero mean, $\int_{M}u=0$.
For a general $u$ write $u=(u-\bar{u})+\bar{u}$, where $\bar{u}=\operatorname{Vol}(M)^{-1}\int_{M}u$. Then
[…]
The first term is controlled by (3). For the second, we note that
$|\bar{u}|^{2}\leq\frac{1}{\operatorname{Vol}(M)}\int_{M}|u|^{2}\leq\frac{1}{v}\int_{M}|u|^{2},$
and hence
$\Bigl(\int_{M}|\bar{u}|^{4}\Bigr)^{1/2}=|\bar{u}|^{2}\,\operatorname{Vol}(M)^{1/2}\leq\frac{\operatorname{Vol}(M)^{1/2}}{v}\int_{M}|u|^{2}\leq\frac{D}{v}\,\frac{1}{D^{2}}\int_{M}|u|^{2}.$
Absorbing all multiplicative constants into $C_{S}(v,D)$ yields (2). ∎
Remark 2.2.
The existence of such a Sobolev constant under a Ricci lower bound, a volume lower bound and a diameter upper bound is classical; see, for example,[15].
$2.$ $L^{2}$ curvature and the curvature equation
As in [8], one may obtain the global $L^{2}$ curvature bound for manifolds with our assumptions. In the case of Einstein metric, the proof is simpler due to the Chern–Gauss–Bonnet formula.
Theorem 2.3 ([8], Theorem 1.13).
There exists $C_{L^{2}}(v,D)>0$ such that
$\int_{M}|\operatorname{Rm}_{g}|^{2}\leq C_{L^{2}}(v,D)\qquad\text{for all }(M^{4},g)\in\mathcal{E}_{1}(4,v,D).$
Proof.
Since $M$ is Einstein, $\operatorname{Ric}_{g}=\lambda g$ with $|\lambda|\leq 3$, so the scalar curvature is constant and equal to $\mathrm{Scal}=4\lambda$. In dimension $4$ the Riemann tensor decomposes as
[…]
where $W$ is the Weyl tensor and $E$ the traceless Ricci tensor. For Einstein metrics we have $E\equiv 0$, hence
(4) $|\operatorname{Rm}|^{2}=|W|^{2}+\frac{\mathrm{Scal}^{2}}{24}.$
The Chern–Gauss–Bonnet formula in dimension $4$ reads
$8\pi^{2}\chi(M)=\int_{M}\Bigl(|W|^{2}-\frac{1}{2}|E|^{2}+\frac{\mathrm{Scal}^{2}}{24}\Bigr)\,d\operatorname{vol}_{g}.$
Again, $E\equiv 0$ for Einstein metrics, so combining this with (4) gives
$\int_{M}|\operatorname{Rm}|^{2}\,d\operatorname{vol}_{g}=8\pi^{2}\chi(M).$
By the Cheeger–Naber finiteness theorem for Einstein 4-manifolds with two-sided Ricci bounds, a lower volume bound and an upper diameter bound (see [8] together with Anderson [2, 3]), the class $\mathcal{E}_{1}(4,v,D)$ contains only finitely many diffeomorphism types. In particular,
$|\chi(M)|\ \leq\ C_{\chi}(v,D)$
for some constant $C_{\chi}(v,D)$ depending only on $v,D$. It follows that
$\int_{M}|\operatorname{Rm}|^{2}\ =\ 8\pi^{2}|\chi(M)|\ \leq\ 8\pi^{2}C_{\chi}(v,D)$
for all $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$, so we may set $C_{L^{2}}(v,D)=8\pi^{2}C_{\chi}(v,D)$. ∎
Next we recall the elliptic equation satisfied by the curvature tensor. For an Einstein metric $\operatorname{Ric}_{g}=\lambda g$, recall the Weitzenböck formula for $\operatorname{Rm}$:
(5) $\Delta\operatorname{Rm}=\operatorname{Rm}*\operatorname{Rm}+\lambda\,\operatorname{Rm},$
where $\Delta$ is the rough Laplacian and $\operatorname{Rm}*\operatorname{Rm}$ denotes a quadratic expression in $\operatorname{Rm}$; see for example [4]. Let
$u:=|\operatorname{Rm}|.$
Then, using the standard Bochner formula and Kato’s inequality $|\nabla u|\leq|\nabla\operatorname{Rm}|$, one obtains a scalar differential inequality of the form
(6) $\Delta u\ \leq\ C_{0}\,(u^{2}+u)$
for some dimensional constant $C_{0}>0$. We briefly justify this.
Indeed, compute
$\frac{1}{2}\Delta u^{2}=\langle\Delta\operatorname{Rm},\operatorname{Rm}\rangle+|\nabla\operatorname{Rm}|^{2}.$
Using (5) and the Cauchy–Schwarz inequality,
$\langle\Delta\operatorname{Rm},\operatorname{Rm}\rangle=\langle\operatorname{Rm}*\operatorname{Rm},\operatorname{Rm}\rangle+\lambda|\operatorname{Rm}|^{2}\geq-C\,|\operatorname{Rm}|^{3}-C\,|\operatorname{Rm}|^{2}$
for some dimensional constant $C>0$. Therefore
$\frac{1}{2}\Delta u^{2}\geq-Cu^{3}-Cu^{2}+|\nabla\operatorname{Rm}|^{2}.$
On the other hand,
$\frac{1}{2}\Delta u^{2}=u\Delta u+|\nabla u|^{2}.$
Combining these identities and using Kato’s inequality $|\nabla u|\leq|\nabla\operatorname{Rm}|$ yields
$u\Delta u+|\nabla u|^{2}\geq-Cu^{3}-Cu^{2}+|\nabla u|^{2},$
hence
$\Delta u\geq-C(u^{2}+u).$
Rewriting this as $\Delta u\leq C_{0}(u^{2}+u)$ with $C_{0}=C$ gives (6).
2.3. $\varepsilon$-regularity via Sobolev and Moser iteration
We now state and prove a standard Moser-iteration lemma adapted to the inequality (6) and the local Sobolev inequality (8).
Restricting (2) to functions supported in a ball $B_{2r}(x)\subset M$ (with $r\leq D$) immediately gives a local inequality of the form
(7) $\Bigl(\int_{B_{2r}(x)}|f|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{B_{2r}(x)}|\nabla f|^{2}+\frac{1}{r^{2}}\int_{B_{2r}(x)}f^{2}\right)^{1/2}$
for all $f\in C_{c}^{\infty}(B_{2r}(x))$, since $1/D^{2}\leq 1/r^{2}$ when $r\leq D$.
Lemma 2.4 (Local $L^{\infty}$ control from small $L^{2}$).
Let $u\geq 0$ satisfy
$\Delta u\leq C_{0}(u^{2}+u)\quad\text{weakly on }B_{2r}(x)\subset M.$
Assume that on $B_{2r}(x)$ an $L^{2}$-Sobolev inequality of the form
(8) $\Bigl(\int_{B_{2r}(x)}|f|^{4}\Bigr)^{1/2}\leq C_{S}^{\prime}\left(\int_{B_{2r}(x)}|\nabla f|^{2}+\frac{1}{r^{2}}\int_{B_{2r}(x)}f^{2}\right)^{1/2}$
holds for all $f\in C_{c}^{\infty}(B_{2r}(x))$, with $C_{S}^{\prime}$ depending only on $C_{S}(v,D)$. Then there exist constants $\varepsilon_{0}=\varepsilon_{0}(C_{0},C_{S}^{\prime})>0$ and $C_{1}=C_{1}(C_{0},C_{S}^{\prime})>0$ such that if
$\int_{B_{2r}(x)}u^{2}\leq\varepsilon_{0},$
then
$\sup_{B_{r}(x)}u\leq C_{1}r^{-2}.$
Proof.
We give the standard Moser iteration argument. Fix a smooth cutoff function $\eta$ with
$0\leq\eta\leq 1,\quad\eta\equiv 1\text{ on }B_{r}(x),\quad\operatorname{supp}\eta\subset B_{2r}(x),\quad|\nabla\eta|\leq\frac{2}{r}.$
Let $p\geq 2$ and consider the test function $\varphi=\eta^{2}u^{p-1}$ in the weak formulation of (6). Multiplying by $\varphi$ and integrating over $B_{2r}(x)$ gives
$\int\langle\nabla u,\nabla\varphi\rangle\leq C_{0}\int\varphi(u^{2}+u),$
where all integrals are over $B_{2r}(x)$ and we suppress the domain in the notation. Expanding the left-hand side,
$\nabla\varphi=2\eta u^{p-1}\nabla\eta+(p-1)\eta^{2}u^{p-2}\nabla u,$
so
$\int\langle\nabla u,\nabla\varphi\rangle=(p-1)\int\eta^{2}u^{p-2}|\nabla u|^{2}+2\int\eta u^{p-1}\langle\nabla u,\nabla\eta\rangle.$
Using Young’s inequality, for any $\delta>0$,
$2\left|\int\eta u^{p-1}\langle\nabla u,\nabla\eta\rangle\right|\leq\delta\int\eta^{2}u^{p-2}|\nabla u|^{2}+\frac{4}{\delta}\int u^{p}|\nabla\eta|^{2}.$
Choose $\delta=\frac{p-1}{2}$ to absorb half of the main term. Then
$\frac{p-1}{2}\int\eta^{2}u^{p-2}|\nabla u|^{2}\leq C_{0}\int\eta^{2}u^{p+1}+C_{0}\int\eta^{2}u^{p}+\frac{8}{p-1}\int u^{p}|\nabla\eta|^{2}.$
Now set $v=\eta u^{p/2}$. Then
$|\nabla v|\leq\eta u^{p/2-1}|\nabla u|+u^{p/2}|\nabla\eta|,$
and one checks that
$\int|\nabla v|^{2}\leq C(p)\int\eta^{2}u^{p-2}|\nabla u|^{2}+C(p)\int u^{p}|\nabla\eta|^{2}$
for some $C(p)$ depending only on $p$. Combining the two inequalities yields
(9) $\int|\nabla v|^{2}\leq C(p)\Bigl(\int\eta^{2}u^{p+1}+\int\eta^{2}u^{p}+\int u^{p}|\nabla\eta|^{2}\Bigr).$
We now apply the local Sobolev inequality (8) to $f=v$:
$\Bigl(\int|v|^{4}\Bigr)^{1/2}\leq C_{S}^{\prime}\left(\int|\nabla v|^{2}+\frac{1}{r^{2}}\int v^{2}\right)^{1/2}.$
Since $v^{2}=\eta^{2}u^{p}\leq u^{p}$ and $|\nabla\eta|\leq 2/r$, (9) and (8) together give
$\Bigl(\int|v|^{4}\Bigr)^{1/2}\leq C_{2}(p,C_{S}^{\prime})\left(\int u^{p+1}+\int u^{p}+\frac{1}{r^{2}}\int u^{p}\right)^{1/2},$
where $C_{2}$ depends only on $p$ and $C_{S}^{\prime}$. Using the smallness of $\int u^{2}$ and Hölder interpolation, one can absorb the $u^{p+1}$ term. More precisely, for $p\geq 2$ we have
$\int u^{p+1}\leq\|u\|_{L^{2}}^{1-\theta}\|u\|_{L^{q}}^{\theta(p+1)}$
with suitable $q>p$ and $\theta\in(0,1)$. Taking $\varepsilon_{0}>0$ sufficiently small so that $\|u\|_{L^{2}(B_{2r})}\leq\varepsilon_{0}$ makes the $u^{p+1}$ term a small perturbation, which can be absorbed into the left-hand side after a standard iteration. The upshot is that there exist exponents $p_{j}$ with $p_{0}=2$ and $p_{j+1}=2p_{j}$ and radii $r_{j}$ with $r\leq r_{j}\leq 2r$, and a constant $C_{3}=C_{3}(C_{0},C_{S}^{\prime})$ such that
$\|u\|_{L^{p_{j+1}}(B_{r_{j+1}})}\leq C_{3}\,r^{-2}\,\|u\|_{L^{p_{j}}(B_{r_{j}})}\quad\text{for all }j\geq 0.$
Iterating this inequality and letting $j\to\infty$ yields
$\|u\|_{L^{\infty}(B_{r}(x))}\leq C_{1}(C_{0},C_{S}^{\prime})\,r^{-2}\,\|u\|_{L^{2}(B_{2r}(x))}.$
If we now assume $\int_{B_{2r}(x)}u^{2}\leq\varepsilon_{0}$ with $\varepsilon_{0}$ fixed, we obtain
$\sup_{B_{r}(x)}u\leq C_{1}r^{-2},$
as claimed. ∎
Combining this with Theorem 2.1 (restricted to $B_{2r}(x)$ as in (7)) and the curvature equation we obtain:
Theorem 2.5 (Einstein $\varepsilon$-regularity via Sobolev).
There exist dimensional constants $\varepsilon_{\mathrm{Ein}}>0$ and $c_{\mathrm{Ein}}>0$, depending only on $C_{S}(v,D)$, such that the following holds.
Let $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ and suppose that $B_{2r}(x)\subset M$ satisfies
$\int_{B_{2r}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}.$
Then
(10) $\sup_{B_{r}(x)}|\operatorname{Rm}|\leq Cr^{-2}$
for some universal constant $C$, and the harmonic radius at every point $y\in B_{r}(x)$ satisfies
$r_{h}(y)\geq c_{\mathrm{Ein}}\,r.$
Proof.
Set $u=|\operatorname{Rm}|$. By (6) we have $\Delta u\leq C_{0}(u^{2}+u)$ on $B_{2r}(x)$. The restriction of the global Sobolev inequality (2) to $B_{2r}(x)$ yields a local inequality of the form (8) with $C_{S}^{\prime}=C_{S}(v,D)$, as in (7).
Applying Lemma 2.4, we obtain constants $\varepsilon_{\mathrm{Ein}}=\varepsilon_{0}(C_{0},C_{S}(v,D))>0$ and $C_{1}=C_{1}(C_{0},C_{S}(v,D))$ such that
$\int_{B_{2r}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}\quad\Longrightarrow\quad\sup_{B_{r}(x)}|\operatorname{Rm}|\leq C_{1}r^{-2},$
which is (10).
To obtain the harmonic radius bound, we use standard elliptic regularity in harmonic coordinates. Fix $y\in B_{r}(x)$. Choose a harmonic coordinate chart $\Phi:B_{\rho}(0)\subset\mathbb{R}^{4}\to M$ with $\Phi(0)=y$, where $\rho>0$ is chosen so small that $B_{2\rho}(y)\subset B_{r}(x)$. In harmonic coordinates, the metric components $g_{ij}$ satisfy a quasi-linear elliptic system of the form
$\Delta g_{ij}=Q_{ij}(g,\partial g)-2\operatorname{Ric}_{ij},$
where $Q$ is quadratic in $\partial g$ and $\operatorname{Ric}_{ij}=\lambda g_{ij}$ is bounded. The curvature bound (10) on $B_{r}(x)$ in turn bounds the coefficients of this system. By standard Schauder or $W^{2,p}$ interior estimates (see, e.g., Anderson [2, 3]), one obtains a uniform $C^{1,\alpha}$ bound on $g_{ij}$ in $B_{\rho/2}(0)$, with constants depending only on the curvature bound and $C_{S}(v,D)$. In particular, by rescaling we can choose $\rho$ comparable to $r$ so that
$\|g_{ij}-\delta_{ij}\|_{C^{0}(B_{\rho/2}(0))}+\rho\,\|\partial g_{ij}\|_{C^{0}(B_{\rho/2}(0))}\leq 10^{-3}.$
This implies that the harmonic radius at $y$ satisfies $r_{h}(y)\geq c_{\mathrm{Ein}}\,r$ for some $c_{\mathrm{Ein}}>0$ depending only on $C_{S}(v,D)$ and the dimension. ∎
Corollary 2.6 (Harmonic radius scale).
Define the dimensionless parameter
$\mathbf{A}(v,D):=\frac{C_{L^{2}}(v,D)}{\varepsilon_{\mathrm{Ein}}},$
where $C_{L^{2}}(v,D)$ is as in Theorem 2.3 and $\varepsilon_{\mathrm{Ein}}$ is as in Theorem 2.5. There exists a decreasing function $\Psi:(0,\infty)\to(0,\infty)$ depending only on $C_{S}(v,D)$ and the dimension, such that for each $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ one can cover $M$ by finitely many balls $B_{r_{i}}(x_{i})$ with
$\int_{B_{2r_{i}}(x_{i})}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}},$
and hence
$r_{h}(y)\geq r_{0}(v,D)\,r_{i}\quad\text{for all }y\in B_{r_{i}}(x_{i}),\qquad r_{0}(v,D):=\Psi(\mathbf{A}(v,D)).$
Proof.
Fix $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$. For each $x\in M$, define
$r(x):=\sup\Bigl\{\,0<\rho\leq D:\ \int_{B_{2\rho}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}\,\Bigr\}.$
By continuity of the map $\rho\mapsto\int_{B_{2\rho}(x)}|\operatorname{Rm}|^{2}$ and compactness of $M$, we have $r(x)>0$ for all $x$ and the family $\{B_{r(x)}(x)\}_{x\in M}$ is an open cover of $M$. By compactness, there is a finite subcover $\{B_{r_{i}}(x_{i})\}_{i=1}^{N}$.
On each ball $B_{2r_{i}}(x_{i})$ we have, by construction,
$\int_{B_{2r_{i}}(x_{i})}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}.$
Thus Theorem 2.5 applies (after possibly replacing $r_{i}$ by $r_{i}/2$), and we obtain
$r_{h}(y)\geq c_{\mathrm{Ein}}\,r_{i}\qquad\text{for all }y\in B_{r_{i}}(x_{i}).$
We may now define $\Psi$ by setting
$\Psi(\mathbf{A}(v,D)):=c_{\mathrm{Ein}},$
viewed as a function of the pair $(C_{S}(v,D),\varepsilon_{\mathrm{Ein}})$ and hence of $\mathbf{A}(v,D)=C_{L^{2}}(v,D)/\varepsilon_{\mathrm{Ein}}$. By enlarging $\Psi$ if necessary for small values of $\mathbf{A}$, we can assume that $\Psi$ is decreasing on $(0,\infty)$. Setting $r_{0}(v,D)=\Psi(\mathbf{A}(v,D))$ yields the claimed statement. ∎
Remark 2.7.
In all subsequent estimates we write $r_{0}(v,D)$ for this harmonic-radius scale. Ultimately, all constants will be expressed as functions of $C_{S}(v,D)$ and $\mathbf{A}(v,D)$.
2.4. Bubble–tree decomposition
We state Cheeger–Naber’s structure theorem [8], which will be used to construct a cover on the manifold.
Theorem 2.8 (Bubble–tree decomposition).
There exist integers $N_{E}(v,D)$, $k_{E}(v,D)$ and a decomposition
$M=\mathcal{B}^{1}\cup\bigcup_{j_{2}=1}^{N_{2}}\mathcal{N}^{2}_{j_{2}}\cup\bigcup_{j_{2}=1}^{N_{2}}\mathcal{B}^{2}_{j_{2}}\cup\dots\cup\bigcup_{j_{k}=1}^{N_{k}}\mathcal{N}^{k}_{j_{k}}\cup\bigcup_{j_{k}=1}^{N_{k}}\mathcal{B}^{k}_{j_{k}}$
such that:
  1. (i)
    (Bodies) If $x\in\mathcal{B}^{i}_{j}$ then $r_{h}(x)\geq r_{0}(v,D)\,\operatorname{diam}(\mathcal{B}^{i}_{j}),$ where $r_{h}$ is the harmonic radius and $r_{0}(v,D)$ is as in Corollary 2.6.
  2. (ii)
    (Necks) Each $\mathcal{N}^{i}_{j}$ is diffeomorphic to an annulus in $\mathbb{R}^{4}/\Gamma^{i}_{j}$, with $|\Gamma^{i}_{j}|\leq C_{\Gamma}(\mathbf{A}(v,D))$ for some explicit function $C_{\Gamma}$.
  3. (iii)
    Each $\mathcal{N}^{i}_{j}$ meets exactly two bodies, a parent $\mathcal{B}^{i-1}_{\cdot}$ and a child $\mathcal{B}^{i}_{j}$, and their intersections are diffeomorphic to $\mathbb{R}\times S^{3}/\Gamma^{i}_{j}$.
  4. (iv)
    The constants $N_{i}\leq N_{E}(v,D)$ and $k\leq k_{E}(v,D)$.
This a specialization of Cheeger–Naber’s bubble–tree construction [8, Theorem 8.64] to the Einstein class, using the $\varepsilon$-regularity Theorem 2.5 and the global $L^{2}$-bound of Theorem 2.3. We denote the adjacency relations between bodies and necks as a rooted tree with root $\mathcal{B}^{1}$ and denote by $\mathcal{T}^{i}_{j}$ the subtree with root $\mathcal{B}^{i}_{j}$.
3. Filling functions and stationary varifolds
In this section we recall how an upper bound for the first homological filling function controls the area of a smallest $2$-dimensional stationary integral varifold, and we explain how to obtain such an upper bound in the Einstein setting. The overall strategy follows our previous work [23]: we construct a cover adapted to the bubble–tree decomposition, pass to the nerve and an associated geodesic graph, and then fill $1$-cycles by combining local geometric fillings with a combinatorial filling argument.
The one new ingredient in the present paper is an improved combinatorial filling lemma (Lemma 3.7 below), based on a sharp bound for integer solutions of linear Diophantine systems due to Borosh–Flahive–Rubin–Treybig [5] together with Hadamard’s inequality. For the reader’s convenience we include a complete proof of this lemma. All other steps are essentially the same as those in [23], up to the replacement of the general Ricci bounds by the Einstein $\varepsilon$-regularity input from Section 2, and we therefore only sketch the arguments or omit proofs with precise references to [23].
3.1. Homological filling functions and stationary varifolds
Let $\mathcal{C}_{k}(M;\mathbb{Z})$ be the group of singular Lipschitz $k$-chains with integer coefficients and $\mathcal{Z}_{k}(M;\mathbb{Z})=\ker\partial_{k}$ the group of $k$-cycles. If $C=\sum_{i}a_{i}f_{i}\in\mathcal{C}_{k}(M;\mathbb{Z})$ with $f_{i}:\Delta^{k}\to M$ Lipschitz, we define its $k$-mass by
$\operatorname{mass}_{k}(C)\;=\;\sum_{i}|a_{i}|\int_{\Delta^{k}}f_{i}^{*}(d\operatorname{vol}_{g}).$
Assuming $H_{1}(M;\mathbb{Z})=0$, the first homological filling function is defined for $l\geq 0$ by
$\operatorname{FH}_{1}(l)\;=\;\sup_{\operatorname{mass}_{1}(Z)\leq l}\;\inf_{\partial C=Z}\,\operatorname{mass}_{2}(C),$
where the supremum is taken over all singular Lipschitz $1$-cycles $Z\in\mathcal{Z}_{1}(M;\mathbb{Z})$ with $\operatorname{mass}_{1}(Z)\leq l$.
The connection between $\operatorname{FH}_{1}$ and minimal $2$-varifolds is provided by the following theorem of Nabutovsky–Rotman.
Theorem 3.1 (Nabutovsky–Rotman [19]).
Let $M$ be a closed $n$-dimensional Riemannian manifold with $H_{1}(M;\mathbb{Z})=0$ and diameter $\operatorname{diam}(M)=D$. Let $A_{\min}(M)$ denote the area of a smallest $2$-dimensional stationary integral varifold in $M$. Then
$A_{\min}(M)\;\leq\;\frac{(n+1)!}{2}\,\operatorname{FH}_{1}(2D).$
In particular, for $n=4$,
$A_{\min}(M)\;\leq\;60\,\operatorname{FH}_{1}(2D).$
Thus, in order to bound $A_{\min}(M)$ from above, it suffices to obtain a linear upper bound for $\operatorname{FH}_{1}(l)$ in terms of the geometric data of $(M,g)$. The rest of this section is devoted to establishing such a bound in the Einstein setting.
3.2. A homological filling inequality for Einstein manifolds
We now explain the geometric-combinatorial construction that leads to the Einstein filling inequality. The starting point is the bubble–tree decomposition of Theorem 2.8, which decomposes $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ into finitely many body regions $\mathcal{B}^{i}_{j}$ and neck regions $\mathcal{N}^{i}_{j}$, with quantitative control on harmonic radius in the bodies and on the topology and geometry of the necks.
The idea is to build:
  • (a)
    a cover of $M$ by harmonic balls inside bodies and by “trapezoids” inside necks, with uniformly controlled overlap,
  • (b)
    the nerve $\mathfrak{N}$ of this cover and a canonical map $f:M\to\mathfrak{N}$, and
  • (c)
    a geodesic graph $\Gamma\subset M$ whose vertices correspond to the elements of the cover and whose edges have uniformly bounded length.
A general $1$-cycle $C$ in $M$ will first be homologically replaced by a sum of cycles supported in single bodies, then approximated by a cycle in $\Gamma$, then pushed to a simplicial cycle in $\mathfrak{N}$ and filled combinatorially using Lemma 3.7. Each $2$-simplex in $\mathfrak{N}$ corresponds to a geodesic triangle in $\Gamma$, which can be filled by a $2$-chain of uniformly bounded area. This yields the desired linear filling inequality.
Bodies. For each body $\mathcal{B}^{i}_{j}$, let
$r^{i}_{j}:=\inf_{x\in\mathcal{B}^{i}_{j}}r_{h}(x)\;\geq\;r_{0}(v,D)\,\operatorname{diam}(\mathcal{B}^{i}_{j}),$
where $r_{h}$ is the harmonic radius and $r_{0}(v,D)$ is the scale from Corollary 2.6. Set
$R^{i}_{j}:=\frac{r^{i}_{j}}{64}.$
By volume comparison and the lower bound on $r^{i}_{j}$, one can cover $\mathcal{B}^{i}_{j}$ by at most $N_{1}(v,D)$ balls of radius $R^{i}_{j}$, and we denote a fixed such subcover by $\mathcal{O}(\mathcal{B}^{i}_{j})$.
Necks. For each neck $\mathcal{N}^{i}_{j}\simeq A_{\rho,2\rho}(0)\subset\mathbb{R}^{4}/\Gamma^{i}_{j}$ via a diffeomorphism $\Phi^{i}_{j}$, we cover the spherical factor $S^{3}/\Gamma^{i}_{j}$ by $m_{i,j}\leq N_{2}(v,D)$ convex balls $B_{r_{c}/4}(z_{\ell})$ and define “trapezoids”
$T_{i,j,\ell}=\Phi^{i}_{j}\bigl((2\sqrt{1-\varepsilon}\,\rho,(2-\sqrt{1-\varepsilon})\,\rho)\times B_{r_{c}/4}(z_{\ell})\bigr),$
contained in slightly larger convex regions
$\bar{T}_{i,j,\ell}=\Phi^{i}_{j}\bigl((\rho/2,2\rho)\times B_{r_{c}}(z_{\ell})\bigr),$
where $\varepsilon>0$ and $r_{c}>0$ are chosen (as in [23]) so that the $\bar{T}_{i,j,\ell}$ are contained in the neck and have controlled overlap.
We then define the global cover
$\mathcal{O}:=\Big(\bigcup_{i,j}\mathcal{O}(\mathcal{B}^{i}_{j})\Big)\;\cup\;\Big(\bigcup_{i,j,\ell}T_{i,j,\ell}\Big).$
The number of sets in $\mathcal{O}$ is bounded by $|\mathcal{O}|\leq\tilde{N}_{E}(v,D)$ for some explicit $\tilde{N}_{E}(v,D)$ depending only on the bubble–tree complexity $N_{E}(v,D)$, $k_{E}(v,D)$ and the covering constants $N_{1}(v,D)$, $N_{2}(v,D)$; see also [23, Sec. 4] for the general Ricci-bounded case.
Let $\mathfrak{N}$ be the nerve of $\mathcal{O}$, and let $f:M\to\mathfrak{N}$ denote the canonical map associated to a fixed partition of unity subordinate to $\mathcal{O}$.
Geodesic graph $\Gamma$. We now construct a geodesic graph $\Gamma\subset M$ whose vertices correspond to the centers of the body balls in $\mathcal{O}(\mathcal{B}^{i}_{j})$ and to chosen reference points in the trapezoids $T_{i,j,\ell}$. Whenever two elements of $\mathcal{O}$ intersect, we connect the corresponding vertices in one of the following ways:
  • (i)
    if both sets lie in the same body, we join the centers by a minimizing geodesic inside that body;
  • (ii)
    if both sets are trapezoids in the same neck, we join the reference points by a short curve contained in the union of the two trapezoids;
  • (iii)
    if one set is a body ball and the other is a trapezoid in a neighbouring neck, we concatenate a short geodesic segment inside the ball with a short curve inside the trapezoid.
By construction, each edge of $\Gamma$ has length bounded by
$\leq C(v,D)\,\max\{R^{i}_{j},\rho\},$
for some explicit constant $C(v,D)$ depending only on the geometric data. This construction is identical to that in [23, Sec. 4], with the only difference that the scales $R^{i}_{j}$ and $\rho$ are now tied to the Einstein harmonic-radius scale $r_{0}(v,D)$.
3.3. Local fillings
We next recall the local geometric fillings that will be used to fill the small pieces of a $1$-cycle inside bodies and necks.
Lemma 3.2.
In a harmonic chart $\Phi:B_{r_{h}}(0)\to M$ with
$\|g-\delta\|_{C^{0}}+r_{h}\|\partial g\|_{C^{0}}\leq 10^{-3},$
any curve $\gamma$ satisfies
$\sqrt{1-10^{-3}}\;\operatorname{length}(\gamma)\;\leq\;\operatorname{length}(\Phi\gamma)\;\leq\;\sqrt{1+2\cdot 10^{-3}}\;\operatorname{length}(\gamma),$
and $\Phi(B_{r_{h}}(0))\supset B_{r_{h}/2}(\Phi(0))$.
Lemma 3.3.
Let $r_{h}(x)>0$ and let $C$ be a $1$-cycle whose image is contained in $B_{r_{h}(x)/2}(x)$. Then there exists a $2$-chain $E$ with $\partial E=C$ and
$\operatorname{mass}_{2}(E)\;\leq\;r_{h}(x)\,\operatorname{mass}_{1}(C).$
Proof.
In harmonic coordinates $\Phi:B_{r_{h}}\to M$, the metric is $(1\pm 10^{-3})$-bilipschitz to the Euclidean metric by Lemma 3.2. Coning $\Phi^{-1}(C)$ to 0 in $\mathbb{R}^{4}$ produces a $2$-chain of Euclidean area $\leq\frac{1}{2}r_{h}\,\operatorname{length}(\Phi^{-1}C)$. Pushing forward via $\Phi$ distorts area by at most a fixed factor close to $1$, which we absorb into the constant. ∎
Lemma 3.4 (Neck contraction).
Let $\mathcal{N}\simeq A_{\rho,2\rho}(0)\subset\mathbb{R}^{4}/\Gamma$ be a neck region in the bubble–tree decomposition, and let $C$ be a $1$-cycle with image in $\mathcal{N}$. Then there exist a $1$-cycle $C^{\prime}$ supported in $\mathcal{N}\cap\mathcal{B}$ (the adjacent child body) and a $2$-chain $E$ such that
$\partial E=C-C^{\prime},\ \operatorname{mass}_{1}(C^{\prime})\leq h_{E}(v,D)\,\operatorname{diam}(\mathcal{B}),\ \operatorname{mass}_{2}(E)\leq C(v,D)\,\rho\,\operatorname{mass}_{1}(C),$
where $h_{E}(v,D)$ depends only on an upper bound for $|\Gamma|$ (hence on $\mathbf{A}(v,D)$ and $C_{S}(v,D)$).
Proof.
The argument is exactly as in [23, Sec. 4.2]. Via $\Phi^{-1}$ we view $C$ as a $1$-cycle in the annulus $A_{\rho,2\rho}(0)\subset\mathbb{R}^{4}/\Gamma$. Radial projection onto a sphere $S_{\lambda}/\Gamma$ with $\lambda\in(\rho,2\rho)$ sweeps out a $2$-chain of area $\lesssim\rho\,\operatorname{mass}_{1}(C)$. The projected cycle is a $1$-cycle in $S^{3}/\Gamma$, whose first homology is finite with generators represented by loops of length $\leq 2\pi\lambda$. Replacing the projected cycle by a representative of its homology class supported in a bounded number of such generators yields $C^{\prime}$ with $\operatorname{mass}_{1}(C^{\prime})\leq h_{E}(v,D)\,\lambda$. Pushing forward by $\Phi$ and controlling the distortion inside the annulus completes the proof. ∎
3.4. Approximation by $\Gamma$ and improved combinatorial filling
The next step is to replace a $1$-cycle lying in a body region by a combinatorial cycle in the geodesic graph $\Gamma$.
Lemma 3.5.
Let $C$ be a $1$-cycle supported in a body region $\mathcal{B}$. Then there exists a $1$-cycle $C^{\prime}\subset\Gamma$ and a $2$-chain $E$ such that
$\partial E=C-C^{\prime},\ \operatorname{mass}_{1}(C^{\prime})\leq C_{\Gamma}(v,D)\,\operatorname{mass}_{1}(C),\ \operatorname{mass}_{2}(E)\leq C_{\Gamma}(v,D)\,\operatorname{mass}_{1}(C).$
Proof.
The proof is identical to [23, Lemma 4.5]. One partitions $C$ into short arcs contained in the balls of the cover $\mathcal{O}(\mathcal{B})$, with length $\leq R^{i}_{j}$. Inside each ball one joins the endpoints of an arc to the corresponding center along edges of $\Gamma$. The difference between the original arc and the resulting path in $\Gamma$ lies in a harmonic ball and hence can be filled using Lemma 3.3. Summing over all arcs gives the desired estimates. ∎
The idea of controlling the combinatorial filling via integer solutions of linear systems was introduced in [19]. We are going to show that it is possible to obtain an upper bound of the number of simplicies needed to fill a simplicial cycle which depends linearly on the size of the cycle. Note that this is independent of the dimension of the simplicial complex.
Let $A\in\mathbb{Z}^{\ell\times n}$, $b\in\mathbb{Z}^{\ell}$, and assume $\mathrm{rank}(A)=m\leq\ell$. Set
$M_{A}:=\max_{1\leq i\leq\ell,\;1\leq j\leq n}|a_{ij}|,\qquad M_{b}:=\max_{1\leq i\leq\ell}|b_{i}|.$
Let $Y$ be the maximum of the absolute values of all $m\times m$ minors of the augmented matrix (A|b)(A\,|\,b):
$Y:=\max\bigl\{\,|\det S|:\ S\text{ is an $m\times m$ submatrix of }(A\,|\,b)\,\bigr\}.$
Theorem 3.6.
Assume that the system $Ax=b$ has an integer solution. Then there exists an integer solution $x=(x_{i})\in\mathbb{Z}^{n}$ such that
(11) $\max_{i}|x_{i}|\;\leq\;m^{m/2}\,\max\{\,M_{A}^{m},\;M_{A}^{m-1}M_{b}\,\}\;=\;m^{m/2}M_{A}^{m-1}\max\{M_{A},M_{b}\}.$
In particular, this upper bound depends only on $m$, $M_{A}$ and $M_{b}$ (and is independent of $n$).
Proof.
By [5, Thm. 1], if $Ax=b$ has an integer solution then there exists an integer solution $x$ with
(12) $\max_{i}|x_{i}|\;\leq\;Y,$
where $Y$ is the maximum of the absolute values of the $m\times m$ minors of (A|b)(A\,|\,b) defined above.
We estimate $Y$ using Hadamard’s inequality. Let $S$ be any $m\times m$ submatrix of (A|b)(A\,|\,b) and denote its columns by $s_{1},\dots,s_{m}\in\mathbb{R}^{m}$. Hadamard’s inequality gives
(13) $|\det S|\;\leq\;\prod_{k=1}^{m}\|s_{k}\|_{2}.$
There are two cases.
Case 1: $S$ is formed entirely by columns of $A$. Then every entry of $S$ has absolute value at most $M_{A}$, hence each column has Euclidean norm $\|s_{k}\|_{2}\leq\sqrt{m}\,M_{A}.$ By (13),
$|\det S|\;\leq\;(\sqrt{m}M_{A})^{m}\;=\;m^{m/2}M_{A}^{m}.$
Case 2: $S$ contains the column $b$. Then one column of $S$ (the one coming from $b$) has norm at most $\sqrt{m}\,M_{b}$, while each of the remaining m1m-1 columns has norm at most $\sqrt{m}\,M_{A}$. Thus
$|\det S|\;\leq\;(\sqrt{m}M_{b})\,(\sqrt{m}M_{A})^{m-1}\;=\;m^{m/2}M_{A}^{m-1}M_{b}.$
Taking the maximum over all $S$ in these two cases yields
$Y\;\leq\;m^{m/2}\,\max\{M_{A}^{m},\;M_{A}^{m-1}M_{b}\}.$
Combining this with (12) gives (11). ∎
Lemma 3.7.
Let $X$ be an $n$-dimensional simplicial complex with $n_{0}$ vertices. Let
$c_{k}=\sum_{i}a_{k}^{i}\,\sigma_{i}^{k}$
be a simplicial $k$-boundary in $X$ with integer coefficients. Then there exists a simplicial $(k+1)$-chain
$c_{k+1}=\sum_{i}a_{k+1}^{i}\,\sigma_{i}^{k+1}$
such that $\partial c_{k+1}=c_{k}$ and
(14) $\max_{i}|a_{k+1}^{i}|\;\leq\;\binom{n_{0}}{k+1}^{\frac{1}{2}\binom{n_{0}}{k+1}}\,\max_{i}|a_{k}^{i}|.$
In particular, since the number $n_{k+1}$ of $(k+1)$-simplices satisfies $n_{k+1}\leq\binom{n_{0}}{k+2}$, we also have
(15) $\sum_{i}|a_{k+1}^{i}|\;\leq\;\binom{n_{0}}{k+2}\,\binom{n_{0}}{k+1}^{\frac{1}{2}\binom{n_{0}}{k+1}}\,\max_{i}|a_{k}^{i}|.$
Proof.
If $c_{k}=0$, the conclusion is trivial with $c_{k+1}=0$, so we assume $c_{k}\neq 0$. Let $C_{j}$ denote the group of $j$-chains of $X$ with integer coefficients. We fix the usual basis of $C_{j}$ given by all oriented $j$-simplices $\sigma^{j}_{1},\dots,\sigma^{j}_{n_{j}}$, where $n_{j}$ is the number of $j$-simplices of $X$.
Consider the boundary map
$\partial_{k+1}\colon C_{k+1}\to C_{k}$
and write its matrix with respect to these bases as $A\in\mathbb{Z}^{n_{k}\times n_{k+1}}.$ By construction, each entry of $A$ lies in $\{0,\pm 1\}$, so
$M_{A}:=\max_{i,j}|a_{ij}|=1.$
Write $c_{k}$ in coordinates as $c_{k}=\sum_{i}a_{k}^{i}\sigma_{i}^{k},$ and let $b\in\mathbb{Z}^{n_{k}}$ be the column vector with entries $b_{i}=a_{k}^{i}$. Then the condition $\partial c_{k+1}=c_{k}$ is exactly the linear system
$Ax=b,$
where $x\in\mathbb{Z}^{n_{k+1}}$ is the coordinate vector of $c_{k+1}$. Since $c_{k}$ is a boundary, this system has at least one integer solution. Set
$M_{b}:=\max_{i}|b_{i}|=\max_{i}|a_{k}^{i}|\ (\geq 1).$
Let $m:=\mathrm{rank}(A)\leq n_{k}$. Applying Theorem 3.6 with this $A$ and $b$, and using $M_{A}=1$, we obtain an integer solution $x$ of $Ax=b$ such that
(16) $\max_{i}|x_{i}|\;\leq\;m^{m/2}\,\max\{M_{A},\;M_{b}\}\;=\;m^{m/2}\,\max_{i}|a_{k}^{i}|.$
Define $c_{k+1}:=\sum_{i}x_{i}\sigma_{i}^{k+1}$; then $\partial c_{k+1}=c_{k}$, and the coefficients $a_{k+1}^{i}$ of $c_{k+1}$ are exactly the $x_{i}$.
Next we bound $m$ in terms of the number of vertices $n_{0}$. Any $k$-simplex of $X$ is determined by its $(k+1)$ vertices, so the number of $k$-simplices satisfies
$n_{k}\leq\binom{n_{0}}{k+1}$.
Hence
$m=\mathrm{rank}(A)\leq n_{k}\leq\binom{n_{0}}{k+1}$.
Combining this with (16) yields
$\max_{i}|a_{k+1}^{i}|=\max_{i}|x_{i}|\;\leq\;m^{m/2}\,\max_{i}|a_{k}^{i}|\;\leq\;\binom{n_{0}}{k+1}^{\frac{1}{2}\binom{n_{0}}{k+1}}\max_{i}|a_{k}^{i}|,$
which is exactly (14). Finally, since there are at most $n_{k+1}\leq\binom{n_{0}}{k+2}$ nonzero coefficients $a_{k+1}^{i}$, we also have
$\sum_{i}|a_{k+1}^{i}|\;\leq\;n_{k+1}\,\max_{i}|a_{k+1}^{i}|\;\leq\;\binom{n_{0}}{k+2}\,\binom{n_{0}}{k+1}^{\frac{1}{2}\binom{n_{0}}{k+1}}\,\max_{i}|a_{k}^{i}|,$
which is (15). ∎
In our application we will only use Lemma 3.7 for $k=1$ and for the simplicial complex $X=\mathfrak{N}$ (the nerve of $\mathcal{O}$).
Lemma 3.8 (Filling geodesic triangles).
There exists a constant $K_{9}(v,D)>0$ such that any geodesic triangle formed by three edges of $\Gamma$ bounds a $2$-chain $E$ with
$\operatorname{mass}_{2}(E)\leq K_{9}(v,D).$
Proof.
As in [23, Lemma 4.6]. If all three vertices lie inside a single body, the triangle lies in a harmonic ball $B_{r_{h}/2}$ and can be filled using Lemma 3.3. If they lie inside a single neck, the triangle lies in the union of at most two trapezoids contained in $\bar{T}_{i,j,\ell}$, and can be filled by a cone in the annulus, with area $\lesssim\rho^{2}$. Mixed configurations are reduced to these two cases. All constants can be controlled in terms of $r_{0}(v,D)$ and $D$. ∎
3.5. Decomposition and the Einstein filling inequality
We now put the pieces together. The first step is to decompose a general $1$-cycle according to the bubble–tree structure.
Lemma 3.9 (Decomposition along the tree).
Let $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ and let $C\in\mathcal{Z}_{1}(M;\mathbb{Z})$ be a singular Lipschitz $1$-cycle. Then there exists a decomposition
$C=\sum_{i=1}^{N}C_{i},$
with $N\leq N_{E}(v,D)\,k_{E}(v,D)$, such that each $C_{i}$ is supported in a single body or a single neck and
$\sum_{i=1}^{N}\operatorname{mass}_{1}(C_{i})\;\leq\;C_{0}(v,D)\,\operatorname{mass}_{1}(C).$
Proof.
As in [23, Sec. 4.3], one cuts $C$ whenever it crosses a boundary component $\partial\mathcal{B}\cap\mathcal{N}$ and reassembles the resulting arcs into cycles supported in individual bodies or necks. The number of such regions is bounded by the bubble–tree complexity $N_{E}(v,D)\,k_{E}(v,D)$, and the total length increases by at most a uniform multiplicative factor $C_{0}(v,D)$, coming from bounded overlap of the decomposition and the uniform control on the diameter of bodies and necks. ∎
Corollary 3.10 (Reduction to bodies).
Let $C$ be as in Lemma 3.9. Then $C$ is homologous to a $1$-cycle $C^{\mathrm{body}}$ supported in the union of body regions, i.e. there exists a $2$-chain $E_{0}$ such that
$\partial E_{0}=C-C^{\mathrm{body}},$
and
$\operatorname{mass}_{1}(C^{\mathrm{body}})\leq C_{1}(v,D)\,\operatorname{mass}_{1}(C),\ \operatorname{mass}_{2}(E_{0})\leq C_{2}(v,D)\,\operatorname{mass}_{1}(C)+C_{3}(v,D).$
Proof.
Apply Lemma 3.4 to each neck-supported component $C_{i}$ in the decomposition of Lemma 3.9. This replaces each $C_{i}$ by a cycle supported in the adjacent body at the expense of adding a $2$-chain whose mass is controlled by $C(v,D)\,\rho\,\operatorname{mass}_{1}(C_{i})$. Summing over all necks and using the uniform bounds on the number and geometry of necks yields the stated estimates. ∎
We are now ready to state and prove the main filling inequality in the Einstein class.
Theorem 3.11 (Einstein filling inequality).
There exist $f_{1}^{\mathrm{Ein}}(v,D),f_{2}^{\mathrm{Ein}}(v,D)>0$ such that for any $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ and any $C\in\mathcal{Z}_{1}(M;\mathbb{Z})$ there exists $E\in\mathcal{C}_{2}(M;\mathbb{Z})$ with
$\partial E=C,\ \operatorname{mass}_{2}(E)\leq f_{1}^{\mathrm{Ein}}(v,D)\,\operatorname{mass}_{1}(C)+f_{2}^{\mathrm{Ein}}(v,D).$
Equivalently, the first homological filling function satisfies
$\operatorname{FH}_{1}(l)\leq f_{1}^{\mathrm{Ein}}(v,D)\,l+f_{2}^{\mathrm{Ein}}(v,D)\ \text{for all }l\geq 0.$
Proof.
By Corollary 3.10, $C$ is homologous to a cycle $C^{\mathrm{body}}$ supported in the union of bodies, with
$\partial E_{0}=C-C^{\mathrm{body}},\ \operatorname{mass}_{2}(E_{0})\leq C_{2}(v,D)\,\operatorname{mass}_{1}(C)+C_{3}(v,D),$
and $\operatorname{mass}_{1}(C^{\mathrm{body}})\leq C_{1}(v,D)\,\operatorname{mass}_{1}(C).$
Next, approximate $C^{\mathrm{body}}$ by a cycle in the geodesic graph $\Gamma$ using Lemma 3.5. We obtain a cycle $\widehat{C}\subset\Gamma$ and a $2$-chain $E_{1}$ such that
$\partial E_{1}=C^{\mathrm{body}}-\widehat{C},\ \operatorname{mass}_{2}(E_{1})\leq C_{\Gamma}(v,D)\,\operatorname{mass}_{1}(C^{\mathrm{body}})\leq C_{\Gamma}(v,D)\,C_{1}(v,D)\,\operatorname{mass}_{1}(C).$
The cycle $\widehat{C}$ maps under $f$ to a simplicial $1$-cycle f(C^)f($\widehat${C}) in the nerve $\mathfrak{N}$. Let $n_{0}$ be the number of vertices of $\mathfrak{N}$, so $n_{0}\leq\tilde{N}_{E}(v,D)$. By construction, the simplicial coefficients of f(C^)f($\widehat${C}) are bounded in absolute value by a constant multiple of the simplicial length $L$ of $\widehat{C}$, and hence by
$\max_{i}|a_{1}^{i}|\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C),$
for some explicit constant $C_{L}(v,D)$.
[…]
$\sum_{i}|a_{2}^{i}|\;\leq\;B_{1}(n_{0})\,\max_{i}|a_{1}^{i}|\;\leq\;B_{1}(\tilde{N}_{E}(v,D))\,C_{L}(v,D)\,\operatorname{mass}_{1}(C),$
where $B_{1}(n_{0})$ is the combinatorial coefficient from (15). Each $2$-simplex in $\mathfrak{N}$ corresponds to a geodesic triangle in $\Gamma$, which bounds a $2$-chain in $M$ of area at most $K_{9}(v,D)$ by Lemma 3.8. Thus $C_{2}$ lifts to a $2$-chain $E_{2}$ in $M$ with
$\partial E_{2}=\widehat{C},\ \operatorname{mass}_{2}(E_{2})\leq K_{9}(v,D)\,\sum_{i}|a_{2}^{i}|\leq K_{9}(v,D)\,B_{1}(\tilde{N}_{E}(v,D))\,C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
Finally, set
$E:=E_{0}+E_{1}+E_{2}.$
Then $\partial E=C$, and combining the three estimates we obtain
$\operatorname{mass}_{2}(E)\leq\bigl(C_{2}+C_{\Gamma}C_{1}+K_{9}B_{1}(\tilde{N}_{E})C_{L}\bigr)\operatorname{mass}_{1}(C)\;+\;C_{3}.$
We may therefore take
$f_{1}^{\mathrm{Ein}}(v,D):=C_{2}+C_{\Gamma}C_{1}+K_{9}\,B_{1}(\tilde{N}_{E})\,C_{L},$
and
$f_{2}^{\mathrm{Ein}}(v,D):=C_{3}(v,D).$
By construction, all constants involved depend only on the harmonic-radius scale $r_{0}(v,D)$, on the Sobolev constant $C_{S}(v,D)$, on upper bounds for the neck groups $|\Gamma^{i}_{j}|$, and on the bubble–tree complexity $\tilde{N}_{E}(v,D)$, $N_{E}(v,D)$, $k_{E}(v,D)$. ∎
Acknowledgements
The author Z. Zhu was supported in part by National Key R&D Program of China 2024YFA1015300 and NSFC 1250010531.

### v2

@@
 $A_{\min}(M,g)\leq F_{\operatorname{Ein}}(v,D),$
-where $F_{\operatorname{Ein}}$ depends only on $(v,D)$ and on quantitative Sobolev and $\varepsilon$–regularity constants for Einstein metrics.
+where $F_{\operatorname{Ein}}$ depends only on $(v,D)$.
+The proof uses the Einstein-specific analytic input of a uniform Sobolev inequality, a global $L^{2}$-curvature bound, and standard $\varepsilon$-regularity/harmonic-radius estimates.
 1. Introduction
@@
 We write $\mathcal{E}_{1}(4,v,D)$ for the class of all closed Einstein $4$-manifolds satisfying (1).
-This paper is a sequel to our previous work [23], where we considered the larger class with $|{\rm Ric}_{g}|\leq 3,\ \operatorname{Vol}(M,g)\geq v>0,\ \operatorname{diam}(M,g)\leq D,\ H_{1}(M;\mathbb{Z})=0,$ and proved a linear upper bound for the first homological filling function $\operatorname{FH}_{1}$ of $(M,g)$.
+This paper is a sequel to our previous work [24], where we considered the larger class with $|{\rm Ric}_{g}|\leq 3,\ \operatorname{Vol}(M,g)\geq v>0,\operatorname{diam}(M,g)\leq D,\ H_{1}(M;\mathbb{Z})=0,$ and proved an affine upper bound for the first homological filling function $\operatorname{FH}_{1}$ of $(M,g)$.
 More precisely, for every $(M^{4},g)\in\mathcal{M}_{1}(4,v,D)$ we showed that there exist functions $f_{1}(v,D),f_{2}(v,D)>0$ such that
@@
 For $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ the curvature solves an elliptic system, and global Sobolev inequalities can be used, following Anderson [2, 3], to derive quantitative $L^{2}$–curvature bounds and $\varepsilon$–regularity with constants depending only on a Sobolev constant.
-This analytic input, combined with the Einstein specialization of the Cheeger–Naber bubble–tree decomposition and an improved combinatorial filling estimate, allows us to obtain an explicit upper bound for $A_{\min}(M,g)$ in terms of $(v,D)$ and the Sobolev constants.
+This analytic input, combined with the Einstein specialization of the Cheeger–Naber bubble–tree decomposition and an improved combinatorial filling estimate, allows us to obtain a quantitative upper bound for $A_{\min}(M,g)$ in terms of $(v,D)$ and the Sobolev constants.
 Our first main result is the following quantitative bound for minimal $2$-varifolds in Einstein $4$-manifolds.
@@
 $A_{\min}(M,g)\ \leq\ F_{\operatorname{Ein}}(v,D).$
-Moreover $F_{\operatorname{Ein}}(v,D)$ can be written explicitly in terms of the Sobolev constant $C_{S}(v,D)$, the global $L^{2}$–curvature bound, and the quantitative Einstein $\varepsilon$–regularity constants.
+Moreover, the dependence of $F_{\operatorname{Ein}}(v,D)$ is reduced to the Sobolev constant $C_{S}(v,D)$, the global $L^{2}$-curvature bound, and the standard Einstein $\varepsilon$-regularity constants.
 As in [24], the proof of Theorem 1.1 proceeds via a homological filling inequality expressed in terms of the first homological filling function $\operatorname{FH}_{1}$ of $(M,g)$.
@@
 Theorem 1.1 is then an immediate consequence of Theorem 1.2 and the min–max theorem of Nabutovsky–Rotman [20], which relates $A_{\min}(M,g)$ to $\operatorname{FH}_{1}$.
-In the Einstein setting we can track all constants arising from the Sobolev inequality, the $L^{2}$–curvature bound and the bubble–tree decomposition, thus making $F_{\operatorname{Ein}}(v,D)$ completely explicit once the analytic data are fixed.
+In the Einstein setting we can track the dependence of the constants arising from the Sobolev inequality, the $L^{2}$-curvature bound and the bubble-tree decomposition, thus reducing the dependence of $F_{\operatorname{Ein}}(v,D)$ to these analytic and combinatorial inputs.
 2. Sobolev inequality and Einstein $\varepsilon$-regularity
@@
 There exists a constant $C_{S}(v,D)>0$ depending only on $v,D$ such that for every $u\in C^{\infty}(M)$ we have
-(2) $\Bigl(\int_{M}|u|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{M}|\nabla u|^{2}+\frac{1}{D^{2}}\int_{M}u^{2}\right)^{1/2}.$
+(2) $\Bigl(\int_{M}|u|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{M}|\nabla u|^{2}+\frac{1}{D^{2}}\int_{M}u^{2}\right).$
 Proof.
-Since $|\operatorname{Ric}_{g}|\leq 3$, we have $\operatorname{Ric}_{g}\geq-3g$. By Bishop–Gromov volume comparison, there exists a dimensional constant $c_{0}>0$ such that
-$\operatorname{Vol}B_{1}(x)\ \geq\ c_{0}\,v\qquad\text{for all }x\in M.$
-In particular $M$ is non-collapsed on the unit scale and has a uniform lower bound on the isoperimetric constant: there exists $I(v,D)>0$ such that for all smooth domains $\Omega\subset M$ with volume at most $\frac{1}{2}\operatorname{Vol}(M)$ one has
-$\operatorname{Vol}(\Omega)^{3/4}\ \leq\ I(v,D)\,\mathcal{H}^{3}(\partial\Omega),$
-where $\mathcal{H}^{3}$ denotes the $3$-dimensional Hausdorff measure induced by $g$.
-By a standard argument (see e.g. Hebey, Nonlinear analysis on manifolds), such inequality is equivalent to an $L^{2}$-Sobolev inequality: there exists $C_{S}(v,D)$ such that
-(3) $\Bigl(\int_{M}|u|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{M}|\nabla u|^{2}\right)^{1/2}$
-for every $u\in C^{\infty}(M)$ with zero mean, $\int_{M}u=0$.
-For a general $u$ write $u=(u-\bar{u})+\bar{u}$, where $\bar{u}=\operatorname{Vol}(M)^{-1}\int_{M}u$. Then
-[…]
-The first term is controlled by (3). For the second, we note that
-$|\bar{u}|^{2}\leq\frac{1}{\operatorname{Vol}(M)}\int_{M}|u|^{2}\leq\frac{1}{v}\int_{M}|u|^{2},$
-and hence
-$\Bigl(\int_{M}|\bar{u}|^{4}\Bigr)^{1/2}=|\bar{u}|^{2}\,\operatorname{Vol}(M)^{1/2}\leq\frac{\operatorname{Vol}(M)^{1/2}}{v}\int_{M}|u|^{2}\leq\frac{D}{v}\,\frac{1}{D^{2}}\int_{M}|u|^{2}.$
-Absorbing all multiplicative constants into $C_{S}(v,D)$ yields (2). ∎
+Under our standing assumptions we have $\operatorname{Ric}_{g}\geq-3g$, $\operatorname{Vol}(M)\geq v$ and $\operatorname{diam}(M)\leq D$.
+A standard consequence is a uniform isoperimetric inequality, hence a uniform Sobolev inequality for mean-zero functions; see, for instance, [16, 2].
+Thus there exists $C^{\prime}_{S}(v,D)$ such that
+$\|u-\bar{u}\|_{L^{4}(M)}^{2}\leq C^{\prime}_{S}(v,D)\,\|\nabla u\|_{L^{2}(M)}^{2},\qquad\bar{u}:=\operatorname{Vol}(M)^{-1}\int_{M}u.$
+In addition, Bishop–Gromov comparison with $\operatorname{Ric}_{g}\geq-3g$ and $\operatorname{diam}(M)\leq D$ gives an upper volume bound $\operatorname{Vol}(M)\leq V_{+}(D)$. Since
+$|\bar{u}|^{2}\leq\operatorname{Vol}(M)^{-1}\|u\|_{L^{2}(M)}^{2}\leq v^{-1}\|u\|_{L^{2}(M)}^{2},$
+we obtain
+$\|\bar{u}\|_{L^{4}(M)}^{2}=|\bar{u}|^{2}\,\operatorname{Vol}(M)^{1/2}\leq v^{-1}V_{+}(D)^{1/2}\,\|u\|_{L^{2}(M)}^{2}\leq C^{\prime\prime}_{S}(v,D)\,\frac{1}{D^{2}}\|u\|_{L^{2}(M)}^{2},$
+after enlarging the constant. Combining the bounds for $u-\bar{u}$ and $\bar{u}$ yields (2). ∎
 Remark 2.2.
@@
 $2.$ $L^{2}$ curvature and the curvature equation
-As in [8], one may obtain the global $L^{2}$ curvature bound for manifolds with our assumptions. In the case of Einstein metric, the proof is simpler due to the Chern–Gauss–Bonnet formula.
-Theorem 2.3 ([8], Theorem 1.13).
+We use the local $L^{2}$ curvature estimate of Cheeger–Naber together with a covering argument to obtain a global bound in our setting.
+Theorem 2.3 ($L^{2}$ curvature bound).
 There exists $C_{L^{2}}(v,D)>0$ such that
@@
 Proof.
-Since $M$ is Einstein, $\operatorname{Ric}_{g}=\lambda g$ with $|\lambda|\leq 3$, so the scalar curvature is constant and equal to $\mathrm{Scal}=4\lambda$.
-In dimension $4$ the Riemann tensor decomposes as
-[…]
-where $W$ is the Weyl tensor and $E$ the traceless Ricci tensor. For Einstein metrics we have $E\equiv 0$, hence
-(4) $|\operatorname{Rm}|^{2}=|W|^{2}+\frac{\mathrm{Scal}^{2}}{24}.$
-The Chern–Gauss–Bonnet formula in dimension $4$ reads
-$8\pi^{2}\chi(M)=\int_{M}\Bigl(|W|^{2}-\frac{1}{2}|E|^{2}+\frac{\mathrm{Scal}^{2}}{24}\Bigr)\,d\operatorname{vol}_{g}.$
-Again, $E\equiv 0$ for Einstein metrics, so combining this with (4) gives
-$\int_{M}|\operatorname{Rm}|^{2}\,d\operatorname{vol}_{g}=8\pi^{2}\chi(M).$
-By the Cheeger–Naber finiteness theorem for Einstein 4-manifolds with two-sided Ricci bounds, a lower volume bound and an upper diameter bound (see [8] together with Anderson [2, 3]), the class $\mathcal{E}_{1}(4,v,D)$ contains only finitely many diffeomorphism types.
-In particular,
-$|\chi(M)|\ \leq\ C_{\chi}(v,D)$
-for some constant $C_{\chi}(v,D)$ depending only on $v,D$. It follows that
-$\int_{M}|\operatorname{Rm}|^{2}\ =\ 8\pi^{2}|\chi(M)|\ \leq\ 8\pi^{2}C_{\chi}(v,D)$
-for all $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$, so we may set $C_{L^{2}}(v,D)=8\pi^{2}C_{\chi}(v,D)$. ∎
+Since $|\operatorname{Ric}_{g}|\leq 3$, Bishop–Gromov comparison and the bounds $\operatorname{Vol}(M)\geq v$, $\operatorname{diam}(M)\leq D$ imply that there exists $v_{1}(v,D)>0$ such that
+$\operatorname{Vol}(B_{1}(x))\geq v_{1}(v,D)\qquad\text{for every }x\in M.$
+Applying the local $L^{2}$ estimate of Cheeger–Naber [8, Theorem 1.13] to each unit ball gives
+$\int_{B_{1}(x)}|\operatorname{Rm}|^{2}\leq C_{\mathrm{CN}}(v_{1}(v,D))\qquad\text{for every }x\in M.$
+Now choose a maximal $\frac{1}{2}$-separated set $\{x_{1},\dots,x_{N}\}\subset M$.
+Then the balls $B_{1}(x_{i})$ cover $M$, while the balls $B_{1/4}(x_{i})$ are disjoint.
+Using again Bishop–Gromov comparison together with the diameter bound, we obtain $N\leq N(v,D)$ for some constant depending only on $(v,D)$.
+Therefore
+$\int_{M}|\operatorname{Rm}|^{2}\leq\sum_{i=1}^{N}\int_{B_{1}(x_{i})}|\operatorname{Rm}|^{2}\leq N(v,D)\,C_{\mathrm{CN}}(v_{1}(v,D)).$
+This gives the desired global bound. ∎
 Next we recall the elliptic equation satisfied by the curvature tensor. For an Einstein metric $\operatorname{Ric}_{g}=\lambda g$, recall the Weitzenböck formula for $\operatorname{Rm}$:
@@
 Rewriting this as $\Delta u\leq C_{0}(u^{2}+u)$ with $C_{0}=C$ gives (4).
-2.3. $\varepsilon$-regularity via Sobolev and Moser iteration
-We now state and prove a standard Moser-iteration lemma adapted to the inequality (6) and the local Sobolev inequality (8).
-Restricting (2) to functions supported in a ball $B_{2r}(x)\subset M$ (with $r\leq D$) immediately gives a local inequality of the form
-(7) $\Bigl(\int_{B_{2r}(x)}|f|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{B_{2r}(x)}|\nabla f|^{2}+\frac{1}{r^{2}}\int_{B_{2r}(x)}f^{2}\right)^{1/2}$
-for all $f\in C_{c}^{\infty}(B_{2r}(x))$, since $1/D^{2}\leq 1/r^{2}$ when $r\leq D$.
-Lemma 2.4 (Local $L^{\infty}$ control from small $L^{2}$).
-Let $u\geq 0$ satisfy
-$\Delta u\leq C_{0}(u^{2}+u)\quad\text{weakly on }B_{2r}(x)\subset M.$
-Assume that on $B_{2r}(x)$ an $L^{2}$-Sobolev inequality of the form
-(8) $\Bigl(\int_{B_{2r}(x)}|f|^{4}\Bigr)^{1/2}\leq C_{S}^{\prime}\left(\int_{B_{2r}(x)}|\nabla f|^{2}+\frac{1}{r^{2}}\int_{B_{2r}(x)}f^{2}\right)^{1/2}$
-holds for all $f\in C_{c}^{\infty}(B_{2r}(x))$, with $C_{S}^{\prime}$ depending only on $C_{S}(v,D)$.
-Then there exist constants $\varepsilon_{0}=\varepsilon_{0}(C_{0},C_{S}^{\prime})>0$ and $C_{1}=C_{1}(C_{0},C_{S}^{\prime})>0$ such that if
-$\int_{B_{2r}(x)}u^{2}\leq\varepsilon_{0},$
+2.3. Einstein $\varepsilon$-regularity
+The next result is standard in dimension four; we record it in the form needed later.
+Theorem 2.4 (Einstein $\varepsilon$-regularity).
+There exist constants $\varepsilon_{\mathrm{Ein}}(v,D)>0$, $C_{\mathrm{Ein}}(v,D)>0$ and $c_{\mathrm{Ein}}(v,D)>0$ such that the following holds.
+Let $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ and suppose that $B_{2r}(x)\subset M$ satisfies
+$\int_{B_{2r}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}(v,D).$
 Then
-$\sup_{B_{r}(x)}u\leq C_{1}r^{-2}.$
+$\sup_{B_{r}(x)}|\operatorname{Rm}|\leq C_{\mathrm{Ein}}(v,D)\,r^{-2},\qquad r_{h}(y)\geq c_{\mathrm{Ein}}(v,D)\,r\quad\text{for all }y\in B_{r}(x).$
 Proof.
-We give the standard Moser iteration argument. Fix a smooth cutoff function $\eta$ with
-$0\leq\eta\leq 1,\quad\eta\equiv 1\text{ on }B_{r}(x),\quad\operatorname{supp}\eta\subset B_{2r}(x),\quad|\nabla\eta|\leq\frac{2}{r}.$
-Let $p\geq 2$ and consider the test function $\varphi=\eta^{2}u^{p-1}$ in the weak formulation of (6). Multiplying by $\varphi$ and integrating over $B_{2r}(x)$ gives
-$\int\langle\nabla u,\nabla\varphi\rangle\leq C_{0}\int\varphi(u^{2}+u),$
-where all integrals are over $B_{2r}(x)$ and we suppress the domain in the notation. Expanding the left-hand side,
-$\nabla\varphi=2\eta u^{p-1}\nabla\eta+(p-1)\eta^{2}u^{p-2}\nabla u,$
-so
-$\int\langle\nabla u,\nabla\varphi\rangle=(p-1)\int\eta^{2}u^{p-2}|\nabla u|^{2}+2\int\eta u^{p-1}\langle\nabla u,\nabla\eta\rangle.$
-Using Young’s inequality, for any $\delta>0$,
-$2\left|\int\eta u^{p-1}\langle\nabla u,\nabla\eta\rangle\right|\leq\delta\int\eta^{2}u^{p-2}|\nabla u|^{2}+\frac{4}{\delta}\int u^{p}|\nabla\eta|^{2}.$
-Choose $\delta=\frac{p-1}{2}$ to absorb half of the main term. Then
-$\frac{p-1}{2}\int\eta^{2}u^{p-2}|\nabla u|^{2}\leq C_{0}\int\eta^{2}u^{p+1}+C_{0}\int\eta^{2}u^{p}+\frac{8}{p-1}\int u^{p}|\nabla\eta|^{2}.$
-Now set $v=\eta u^{p/2}$. Then
-$|\nabla v|\leq\eta u^{p/2-1}|\nabla u|+u^{p/2}|\nabla\eta|,$
-and one checks that
-$\int|\nabla v|^{2}\leq C(p)\int\eta^{2}u^{p-2}|\nabla u|^{2}+C(p)\int u^{p}|\nabla\eta|^{2}$
-for some $C(p)$ depending only on $p$. Combining the two inequalities yields
-(9) $\int|\nabla v|^{2}\leq C(p)\Bigl(\int\eta^{2}u^{p+1}+\int\eta^{2}u^{p}+\int u^{p}|\nabla\eta|^{2}\Bigr).$
-We now apply the local Sobolev inequality (8) to $f=v$:
-$\Bigl(\int|v|^{4}\Bigr)^{1/2}\leq C_{S}^{\prime}\left(\int|\nabla v|^{2}+\frac{1}{r^{2}}\int v^{2}\right)^{1/2}.$
-Since $v^{2}=\eta^{2}u^{p}\leq u^{p}$ and $|\nabla\eta|\leq 2/r$, (9) and (8) together give
-$\Bigl(\int|v|^{4}\Bigr)^{1/2}\leq C_{2}(p,C_{S}^{\prime})\left(\int u^{p+1}+\int u^{p}+\frac{1}{r^{2}}\int u^{p}\right)^{1/2},$
-where $C_{2}$ depends only on $p$ and $C_{S}^{\prime}$. Using the smallness of $\int u^{2}$ and Hölder interpolation, one can absorb the $u^{p+1}$ term. More precisely, for $p\geq 2$ we have
-$\int u^{p+1}\leq\|u\|_{L^{2}}^{1-\theta}\|u\|_{L^{q}}^{\theta(p+1)}$
-with suitable $q>p$ and $\theta\in(0,1)$.
-Taking $\varepsilon_{0}>0$ sufficiently small so that $\|u\|_{L^{2}(B_{2r})}\leq\varepsilon_{0}$ makes the $u^{p+1}$ term a small perturbation, which can be absorbed into the left-hand side after a standard iteration.
-The upshot is that there exist exponents $p_{j}$ with $p_{0}=2$ and $p_{j+1}=2p_{j}$ and radii $r_{j}$ with $r\leq r_{j}\leq 2r$, and a constant $C_{3}=C_{3}(C_{0},C_{S}^{\prime})$ such that
-$\|u\|_{L^{p_{j+1}}(B_{r_{j+1}})}\leq C_{3}\,r^{-2}\,\|u\|_{L^{p_{j}}(B_{r_{j}})}\quad\text{for all }j\geq 0.$
-Iterating this inequality and letting $j\to\infty$ yields
-$\|u\|_{L^{\infty}(B_{r}(x))}\leq C_{1}(C_{0},C_{S}^{\prime})\,r^{-2}\,\|u\|_{L^{2}(B_{2r}(x))}.$
-If we now assume $\int_{B_{2r}(x)}u^{2}\leq\varepsilon_{0}$ with $\varepsilon_{0}$ fixed, we obtain
-$\sup_{B_{r}(x)}u\leq C_{1}r^{-2},$
-as claimed. ∎
-Combining this with Theorem 2.1 (restricted to $B_{2r}(x)$ as in (7)) and the curvature equation we obtain:
-Theorem 2.5 (Einstein $\varepsilon$-regularity via Sobolev).
-There exist dimensional constants $\varepsilon_{\mathrm{Ein}}>0$ and $c_{\mathrm{Ein}}>0$, depending only on $C_{S}(v,D)$, such that the following holds.
-Let $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ and suppose that $B_{2r}(x)\subset M$ satisfies
-$\int_{B_{2r}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}.$
-Then
-(10) $\sup_{B_{r}(x)}|\operatorname{Rm}|\leq Cr^{-2}$
-for some universal constant $C$, and the harmonic radius at every point $y\in B_{r}(x)$ satisfies
-$r_{h}(y)\geq c_{\mathrm{Ein}}\,r.$
-Proof.
-Set $u=|\operatorname{Rm}|$.
-By (6) we have $\Delta u\leq C_{0}(u^{2}+u)$ on $B_{2r}(x)$.
-The restriction of the global Sobolev inequality (2) to $B_{2r}(x)$ yields a local inequality of the form (8) with $C_{S}^{\prime}=C_{S}(v,D)$, as in (7).
-Applying Lemma 2.4, we obtain constants $\varepsilon_{\mathrm{Ein}}=\varepsilon_{0}(C_{0},C_{S}(v,D))>0$ and $C_{1}=C_{1}(C_{0},C_{S}(v,D))$ such that
-$\int_{B_{2r}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}\quad\Longrightarrow\quad\sup_{B_{r}(x)}|\operatorname{Rm}|\leq C_{1}r^{-2},$
-which is (10).
-To obtain the harmonic radius bound, we use standard elliptic regularity in harmonic coordinates.
-Fix $y\in B_{r}(x)$.
-Choose a harmonic coordinate chart $\Phi:B_{\rho}(0)\subset\mathbb{R}^{4}\to M$ with $\Phi(0)=y$, where $\rho>0$ is chosen so small that $B_{2\rho}(y)\subset B_{r}(x)$.
-In harmonic coordinates, the metric components $g_{ij}$ satisfy a quasi-linear elliptic system of the form
-$\Delta g_{ij}=Q_{ij}(g,\partial g)-2\operatorname{Ric}_{ij},$
-where $Q$ is quadratic in $\partial g$ and $\operatorname{Ric}_{ij}=\lambda g_{ij}$ is bounded.
-The curvature bound (10) on $B_{r}(x)$ in turn bounds the coefficients of this system.
-By standard Schauder or $W^{2,p}$ interior estimates (see, e.g., Anderson [2, 3]), one obtains a uniform $C^{1,\alpha}$ bound on $g_{ij}$ in $B_{\rho/2}(0)$, with constants depending only on the curvature bound and $C_{S}(v,D)$.
-In particular, by rescaling we can choose $\rho$ comparable to $r$ so that
-$\|g_{ij}-\delta_{ij}\|_{C^{0}(B_{\rho/2}(0))}+\rho\,\|\partial g_{ij}\|_{C^{0}(B_{\rho/2}(0))}\leq 10^{-3}.$
-This implies that the harmonic radius at $y$ satisfies $r_{h}(y)\geq c_{\mathrm{Ein}}\,r$ for some $c_{\mathrm{Ein}}>0$ depending only on $C_{S}(v,D)$ and the dimension. ∎
-Corollary 2.6 (Harmonic radius scale).
-Define the dimensionless parameter
-$\mathbf{A}(v,D):=\frac{C_{L^{2}}(v,D)}{\varepsilon_{\mathrm{Ein}}},$
-where $C_{L^{2}}(v,D)$ is as in Theorem 2.3 and $\varepsilon_{\mathrm{Ein}}$ is as in Theorem 2.5.
-There exists a decreasing function $\Psi:(0,\infty)\to(0,\infty)$ depending only on $C_{S}(v,D)$ and the dimension, such that for each $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ one can cover $M$ by finitely many balls $B_{r_{i}}(x_{i})$ with
-$\int_{B_{2r_{i}}(x_{i})}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}},$
-and hence
-$r_{h}(y)\geq r_{0}(v,D)\,r_{i}\quad\text{for all }y\in B_{r_{i}}(x_{i}),\qquad r_{0}(v,D):=\Psi(\mathbf{A}(v,D)).$
-Proof.
-Fix $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$. For each $x\in M$, define
-$r(x):=\sup\Bigl\{\,0<\rho\leq D:\ \int_{B_{2\rho}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}\,\Bigr\}.$
-By continuity of the map $\rho\mapsto\int_{B_{2\rho}(x)}|\operatorname{Rm}|^{2}$ and compactness of $M$, we have $r(x)>0$ for all $x$ and the family $\{B_{r(x)}(x)\}_{x\in M}$ is an open cover of $M$.
-By compactness, there is a finite subcover $\{B_{r_{i}}(x_{i})\}_{i=1}^{N}$.
-On each ball $B_{2r_{i}}(x_{i})$ we have, by construction,
-$\int_{B_{2r_{i}}(x_{i})}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}.$
-Thus Theorem 2.5 applies (after possibly replacing $r_{i}$ by $r_{i}/2$), and we obtain
-$r_{h}(y)\geq c_{\mathrm{Ein}}\,r_{i}\qquad\text{for all }y\in B_{r_{i}}(x_{i}).$
-We may now define $\Psi$ by setting
-$\Psi(\mathbf{A}(v,D)):=c_{\mathrm{Ein}},$
-viewed as a function of the pair $(C_{S}(v,D),\varepsilon_{\mathrm{Ein}})$ and hence of $\mathbf{A}(v,D)=C_{L^{2}}(v,D)/\varepsilon_{\mathrm{Ein}}$.
-By enlarging $\Psi$ if necessary for small values of $\mathbf{A}$, we can assume that $\Psi$ is decreasing on $(0,\infty)$.
-Setting $r_{0}(v,D)=\Psi(\mathbf{A}(v,D))$ yields the claimed statement.
+After rescaling we may assume $r=1$.
+The hypotheses place us in the standard noncollapsed Einstein setting: the Ricci tensor is uniformly bounded, local volume ratios are bounded below in terms of $(v,D)$, and the Sobolev constant is controlled by Theorem 2.1.
+The usual $\varepsilon$-regularity theorem for Einstein $4$-manifolds therefore applies and yields the pointwise curvature bound together with the harmonic-radius estimate on $B_{1}(x)$; see Anderson [2, 3], Cheeger–Tian [9], and Cheeger–Naber [8].
+Rescaling back proves the statement.
 ∎
 Remark 2.5.
-In all subsequent estimates we write $r_{0}(v,D)$ for this harmonic-radius scale. Ultimately, all constants will be expressed as functions of $C_{S}(v,D)$ and $\mathbf{A}(v,D)$.
+In what follows we only use the existence of the constants in Theorems 2.1, 2.3 and 2.4.
+The global harmonic-radius scale entering the bubble–tree decomposition will be taken directly from Cheeger–Naber’s theorem.
 2.4. Bubble–tree decomposition
@@
 Theorem 2.6 (Bubble–tree decomposition).
-There exist integers $N_{E}(v,D)$, $k_{E}(v,D)$ and a decomposition
+There exist constants $r_{0}(v,D)>0$, $C_{\Gamma}(v,D)>0$ and integers $N_{E}(v,D)$, $k_{E}(v,D)$ together with a decomposition
 $M=\mathcal{B}^{1}\cup\bigcup_{j_{2}=1}^{N_{2}}\mathcal{N}^{2}_{j_{2}}\cup\bigcup_{j_{2}=1}^{N_{2}}\mathcal{B}^{2}_{j_{2}}\cup\dots\cup\bigcup_{j_{k}=1}^{N_{k}}\mathcal{N}^{k}_{j_{k}}\cup\bigcup_{j_{k}=1}^{N_{k}}\mathcal{B}^{k}_{j_{k}}$
@@
   1. (i)
-    (Bodies) If $x\in\mathcal{B}^{i}_{j}$ then $r_{h}(x)\geq r_{0}(v,D)\,\operatorname{diam}(\mathcal{B}^{i}_{j}),$ where $r_{h}$ is the harmonic radius and $r_{0}(v,D)$ is as in Corollary 2.6.
+    (Bodies) If $x\in\mathcal{B}^{i}_{j}$ then
+    $r_{h}(x)\geq r_{0}(v,D)\,\operatorname{diam}(\mathcal{B}^{i}_{j}).$
   2. (ii)
-    (Necks) Each $\mathcal{N}^{i}_{j}$ is diffeomorphic to an annulus in $\mathbb{R}^{4}/\Gamma^{i}_{j}$, with $|\Gamma^{i}_{j}|\leq C_{\Gamma}(\mathbf{A}(v,D))$ for some explicit function $C_{\Gamma}$.
+    (Necks) Each $\mathcal{N}^{i}_{j}$ is diffeomorphic to an annulus in $\mathbb{R}^{4}/\Gamma^{i}_{j}$, with $|\Gamma^{i}_{j}|\leq C_{\Gamma}(v,D)$.
   3. (iii)
@@
     The constants $N_{i}\leq N_{E}(v,D)$ and $k\leq k_{E}(v,D)$.
-This a specialization of Cheeger–Naber’s bubble–tree construction [8, Theorem 8.64] to the Einstein class, using the $\varepsilon$-regularity Theorem 2.5 and the global $L^{2}$-bound of Theorem 2.3.
+This is the specialization of Cheeger–Naber’s bubble–tree construction [8, Theorem 8.63] to the Einstein class.
+The displayed decomposition above is their formula (8.64).
 We denote the adjacency relations between bodies and necks as a rooted tree with root $\mathcal{B}^{1}$ and denote by $\mathcal{T}^{i}_{j}$ the subtree with root $\mathcal{B}^{i}_{j}$.
@@
 $A_{\min}(M)\;\leq\;60\,\operatorname{FH}_{1}(2D).$
-Thus, in order to bound $A_{\min}(M)$ from above, it suffices to obtain a linear upper bound for $\operatorname{FH}_{1}(l)$ in terms of the geometric data of $(M,g)$.
+Thus, in order to bound $A_{\min}(M)$ from above, it suffices to obtain a affine upper bound for $\operatorname{FH}_{1}(l)$ in terms of the geometric data of $(M,g)$.
 The rest of this section is devoted to establishing such a bound in the Einstein setting.
@@
     a geodesic graph $\Gamma\subset M$ whose vertices correspond to the elements of the cover and whose edges have uniformly bounded length.
-A general $1$-cycle $C$ in $M$ will first be homologically replaced by a sum of cycles supported in single bodies, then approximated by a cycle in $\Gamma$, then pushed to a simplicial cycle in $\mathfrak{N}$ and filled combinatorially using Lemma 3.7.
+A general $1$-cycle $C$ in $M$ will first be homologically replaced by a sum of cycles supported in single bodies, then approximated by a cycle in $\Gamma$, then straightened via the standard graph-to-nerve procedure of [24, Sec.
+4] to a simplicial cycle in $\mathfrak{N}$ and filled combinatorially using Lemma 3.7.
 Each $2$-simplex in $\mathfrak{N}$ corresponds to a geodesic triangle in $\Gamma$, which can be filled by a $2$-chain of uniformly bounded area.
-This yields the desired linear filling inequality.
+This yields the desired affine filling inequality.
 Bodies. For each body $\mathcal{B}^{i}_{j}$, let
 $r^{i}_{j}:=\inf_{x\in\mathcal{B}^{i}_{j}}r_{h}(x)\;\geq\;r_{0}(v,D)\,\operatorname{diam}(\mathcal{B}^{i}_{j}),$
-where $r_{h}$ is the harmonic radius and $r_{0}(v,D)$ is the scale from Corollary 2.6. Set
+where $r_{h}$ is the harmonic radius and $r_{0}(v,D)$ is the constant from Theorem 2.6. Set
 $R^{i}_{j}:=\frac{r^{i}_{j}}{64}.$
@@
 $\mathcal{O}:=\Big(\bigcup_{i,j}\mathcal{O}(\mathcal{B}^{i}_{j})\Big)\;\cup\;\Big(\bigcup_{i,j,\ell}T_{i,j,\ell}\Big).$
-The number of sets in $\mathcal{O}$ is bounded by $|\mathcal{O}|\leq\tilde{N}_{E}(v,D)$ for some explicit $\tilde{N}_{E}(v,D)$ depending only on the bubble–tree complexity $N_{E}(v,D)$, $k_{E}(v,D)$ and the covering constants $N_{1}(v,D)$, $N_{2}(v,D)$; see also [23, Sec.
+The number of sets in $\mathcal{O}$ is bounded by $|\mathcal{O}|\leq\tilde{N}_{E}(v,D)$ for some $\tilde{N}_{E}(v,D)$ depending only on the bubble–tree complexity $N_{E}(v,D)$, $k_{E}(v,D)$ and the covering constants $N_{1}(v,D)$, $N_{2}(v,D)$; see also [24, Sec.
 4] for the general Ricci-bounded case.
@@
 $\leq C(v,D)\,\max\{R^{i}_{j},\rho\},$
-for some explicit constant $C(v,D)$ depending only on the geometric data.
-This construction is identical to that in [23, Sec.
-4], with the only difference that the scales $R^{i}_{j}$ and $\rho$ are now tied to the Einstein harmonic-radius scale $r_{0}(v,D)$.
+for some constant $C(v,D)$ depending only on the geometric data.
+This construction is identical to that in [24, Sec.
+4], with the only difference that the scales $R^{i}_{j}$ and $\rho$ are now tied to the harmonic-radius scale $r_{0}(v,D)$ furnished by Theorem 2.6.
 3.3. Local fillings
@@
 $\partial E=C-C^{\prime},\ \operatorname{mass}_{1}(C^{\prime})\leq h_{E}(v,D)\,\operatorname{diam}(\mathcal{B}),\ \operatorname{mass}_{2}(E)\leq C(v,D)\,\rho\,\operatorname{mass}_{1}(C),$
-where $h_{E}(v,D)$ depends only on an upper bound for $|\Gamma|$ (hence on $\mathbf{A}(v,D)$ and $C_{S}(v,D)$).
+where $h_{E}(v,D)$ depends only on the uniform bound $C_{\Gamma}(v,D)$ from Theorem 2.6.
 Proof.
-The argument is exactly as in [23, Sec.
+The argument is exactly as in [24, Sec.
 4.2].
@@
 Proof.
-The proof is identical to [23, Lemma 4.5].
+The proof is identical to [24, Lemma 4.5].
 One partitions $C$ into short arcs contained in the balls of the cover $\mathcal{O}(\mathcal{B})$, with length $\leq R^{i}_{j}$.
@@
 The idea of controlling the combinatorial filling via integer solutions of linear systems was introduced in [20].
-We are going to show that it is possible to obtain an upper bound of the number of simplicies needed to fill a simplicial cycle which depends linearly on the size of the cycle.
+We show that one can obtain an upper bound for the number of simplices needed to fill a simplicial cycle which depends linearly on the size of the cycle.
 Note that this is independent of the dimension of the simplicial complex.
@@
 Proof.
-As in [23, Lemma 4.6].
+As in [24, Lemma 4.6].
 If all three vertices lie inside a single body, the triangle lies in a harmonic ball $B_{r_{h}/2}$ and can be filled using Lemma 3.3.
@@
 Proof.
-As in [23, Sec.
+As in [24, Sec.
 4.3], one cuts $C$ whenever it crosses a boundary component $\partial\mathcal{B}\cap\mathcal{N}$ and reassembles the resulting arcs into cycles supported in individual bodies or necks.
@@
 By Corollary 3.10, $C$ is homologous to a cycle $C^{\mathrm{body}}$ supported in the union of bodies, with
-$\partial E_{0}=C-C^{\mathrm{body}},\ \operatorname{mass}_{2}(E_{0})\leq C_{2}(v,D)\,\operatorname{mass}_{1}(C)+C_{3}(v,D),$
+$\partial E_{0}=C-C^{\mathrm{body}},\quad\operatorname{mass}_{2}(E_{0})\leq C_{2}(v,D)\,\operatorname{mass}_{1}(C)+C_{3}(v,D),$
 and $\operatorname{mass}_{1}(C^{\mathrm{body}})\leq C_{1}(v,D)\,\operatorname{mass}_{1}(C).$
@@
 $\partial E_{1}=C^{\mathrm{body}}-\widehat{C},\ \operatorname{mass}_{2}(E_{1})\leq C_{\Gamma}(v,D)\,\operatorname{mass}_{1}(C^{\mathrm{body}})\leq C_{\Gamma}(v,D)\,C_{1}(v,D)\,\operatorname{mass}_{1}(C).$
-The cycle $\widehat{C}$ maps under $f$ to a simplicial $1$-cycle f(C^)f($\widehat${C}) in the nerve $\mathfrak{N}$.
+The remaining step is the chain-level passage from the graph $\Gamma$ to the nerve $\mathfrak{N}$.
+Instead of using the partition-of-unity map directly on $\widehat{C}$, we invoke the graph-to-nerve straightening procedure from [24, Sec.
+4].
+Applied to the graph cycle $\widehat{C}$, it yields a simplicial $1$-boundary
+$c_{1}=\sum_{i}a_{1}^{i}\,\sigma_{i}^{1}$
+in $\mathfrak{N}$ and a $2$-chain $E_{\mathrm{st}}$ in $M$ such that
+$\partial E_{\mathrm{st}}=\widehat{C}-\gamma(c_{1}),$
+where $\gamma(c_{1})$ denotes the edge cycle in $\Gamma$ corresponding to $c_{1}$, and
+$\sum_{i}|a_{1}^{i}|\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C),\qquad\operatorname{mass}_{2}(E_{\mathrm{st}})\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
+In particular,
+$\max_{i}|a_{1}^{i}|\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
 Let $n_{0}$ be the number of vertices of $\mathfrak{N}$, so $n_{0}\leq\tilde{N}_{E}(v,D)$.
-By construction, the simplicial coefficients of f(C^)f($\widehat${C}) are bounded in absolute value by a constant multiple of the simplicial length $L$ of $\widehat{C}$, and hence by
-$\max_{i}|a_{1}^{i}|\leq C_{L}(v,D)\,\operatorname{mass}_{1}(C),$
-for some explicit constant $C_{L}(v,D)$.
-[…]
+Since $c_{1}$ is a simplicial boundary in $\mathfrak{N}$, Lemma 3.7 with $k=1$ yields a simplicial $2$-chain $C_{2}=\sum a_{2}^{i}\sigma_{i}^{2}$ with $\partial C_{2}=c_{1}$ and
 $\sum_{i}|a_{2}^{i}|\;\leq\;B_{1}(n_{0})\,\max_{i}|a_{1}^{i}|\;\leq\;B_{1}(\tilde{N}_{E}(v,D))\,C_{L}(v,D)\,\operatorname{mass}_{1}(C),$
@@
 Thus $C_{2}$ lifts to a $2$-chain $E_{2}$ in $M$ with
-$\partial E_{2}=\widehat{C},\ \operatorname{mass}_{2}(E_{2})\leq K_{9}(v,D)\,\sum_{i}|a_{2}^{i}|\leq K_{9}(v,D)\,B_{1}(\tilde{N}_{E}(v,D))\,C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
+$\partial E_{2}=\gamma(c_{1}),\ \operatorname{mass}_{2}(E_{2})\leq K_{9}(v,D)\,\sum_{i}|a_{2}^{i}|\leq K_{9}(v,D)\,B_{1}(\tilde{N}_{E}(v,D))\,C_{L}(v,D)\,\operatorname{mass}_{1}(C).$
 Finally, set
-$E:=E_{0}+E_{1}+E_{2}.$
+$E:=E_{0}+E_{1}+E_{\mathrm{st}}+E_{2}.$
 Then $\partial E=C$, and combining the three estimates we obtain
-$\operatorname{mass}_{2}(E)\leq\bigl(C_{2}+C_{\Gamma}C_{1}+K_{9}B_{1}(\tilde{N}_{E})C_{L}\bigr)\operatorname{mass}_{1}(C)\;+\;C_{3}.$
+$\operatorname{mass}_{2}(E)\leq\bigl(C_{2}+C_{\Gamma}C_{1}+C_{L}+K_{9}B_{1}(\tilde{N}_{E})C_{L}\bigr)\operatorname{mass}_{1}(C)\;+\;C_{3}.$
 We may therefore take
-$f_{1}^{\mathrm{Ein}}(v,D):=C_{2}+C_{\Gamma}C_{1}+K_{9}\,B_{1}(\tilde{N}_{E})\,C_{L},$
+$f_{1}^{\mathrm{Ein}}(v,D):=C_{2}+C_{\Gamma}C_{1}+C_{L}+K_{9}\,B_{1}(\tilde{N}_{E})\,C_{L},$
 and
 $f_{2}^{\mathrm{Ein}}(v,D):=C_{3}(v,D).$
-By construction, all constants involved depend only on the harmonic-radius scale $r_{0}(v,D)$, on the Sobolev constant $C_{S}(v,D)$, on upper bounds for the neck groups $|\Gamma^{i}_{j}|$, and on the bubble–tree complexity $\tilde{N}_{E}(v,D)$, $N_{E}(v,D)$, $k_{E}(v,D)$.
-∎
+[…]
 Acknowledgements